# Inv-Entropy: A Fully Probabilistic Framework for Uncertainty Quantification in Language Models

### v1

@@ 2.4 Inv-Entropy via bootstrapping and Monte Carlo
A natural measure of uncertainty is the conditional sample entropy $H(X_{n}\mid Y_{n})$, which quantifies the ambiguity of the input set $X_{n}$ given the corresponding output set $Y_{n}$:
$H(X_{n}\mid Y_{n})\triangleq-\operatorname{trace}(\text{P}_{y}\text{P}_{x}\odot\log(\text{P}_{y}\text{P}_{x}))=-\sum_{i=0}^{n}P(x_{i}\mid y_{i})\log P(x_{i}\mid y_{i}),$ (5)
where $\odot$ denotes a Hadamard product. This quantity reflects the alignment between $\text{P}_{y}$ and $\text{P}_{x}$. High entropy indicates that semantically similar inputs lead to different output or semantically different outputs lead to similar input, revealing uncertainty in the model’s behavior.
That said, it is important to realize that $X_{n}$ does not map to a unique $Y_{n}$ due to the inherent stochasticity of $f$, as LLMs can produce different outputs for the same input. Hence, beyond perturbations that capture epistemic uncertainty, replications are needed to capture aleatoric uncertainty from sampling randomness. To this end, we replicate multiple responses per question: each perturbation $S_{x_{i}}\in\mathcal{P}er(S_{x_{0}})$ is passed to the LLM $r$ times, generating a replicated set $\mathcal{R}_{i}=\{y_{i,1},\dots,y_{i,r}\}$, where $y_{i,\cdot}=\psi(f(S_{x_{i}}))$.
@@
With this, we are now able to generate $S$ bootstrapped samples $Y_{n}^{(s)}=\{y_{0}^{(s)},\dots,y_{n}^{(s)}\}$ for $s\in\{1,\dots,S\}$ corresponding to $X_{n}$, where $y_{i}^{(s)}\sim\mathcal{R}_{i}$ for $i\in\{0,\dots,n\}$ is sampled with replacement. Each bootstrapped output embedding set $Y_{n}^{(s)}$, together with the input embedding set $X_{n}$, induces a probabilistic framework that yields an instance of our UQ measure, Inv-Entropy. Accordingly, we apply Monte Carlo estimation to compute the average Inv-Entropy across these instances, resulting in the final UQ measure shown below:
$\widehat{H}(X\mid Y)=\frac{1}{S}\sum_{s=1}^{S}H(X_{n}\mid Y_{n}^{(s)}).$ (6)
@@
3:LLM Generation: Input each element in $\mathcal{P}er(S_{x_{0}})$, $r$ times into $f$ and obtain $\{\mathcal{R}_{i}\}_{i=0}^{n}$.
4:for $s=1$ to $S$ do
5:     Sample $Y_{n}^{(s)}=\{y_{0}^{(s)},\dots,y_{n}^{(s)}\}$
6:     Compute $\text{P}_{y}^{(s)}$ using (2)
7:     Compute $P^{(s)}(x_{i}|y_{i})=\text{P}^{(s)}_{y}\cdot\text{P}_{x}[i,i]$ (from (4)) for $\forall$ $i$
8:     Compute $H(X_{n}\mid Y_{n}^{(s)})$ using (5)
@@ 2.5 GAAP
In this section, we present a genetic algorithm-based adversarial perturbation (GAAP) that progressively modifies the semantic input $S_{x_{0}}$ to generate controlled perturbations, $\mathcal{P}er(S_{x_{0}})$. Below we highlight our overarching framework, while algorithmic details are relegated to Appendix C. As shown in Fig. 3, the process consists of an initialization step and multiple iterative procedures. In the initialization step, we construct a population $\text{Pop}_{0}(S_{x_{0}})$ consisting of perturbed versions of $S_{x_{0}}$. More specifically, each text in $\text{Pop}_{0}(S_{x_{0}})$ is derived from $S_{x_{0}}$ by replacing one keyword with a synonym, hypernym, hyponym from WordNet (Miller, 1995), or a deletion.
Figure 3: Illustration of GAAP on a TriviaQA (Joshi et al., 2017) question.
@@ 3 Experiments
AUROC ($\uparrow$) Semantic Entropy 0.579 $\pm$ 0.044 0.679 $\pm$ 0.045 0.661 $\pm$ 0.048 0.518 $\pm$ 0.048 0.589 $\pm$ 0.052
VU 0.695 $\pm$ 0.060 0.480 $\pm$ 0.060 0.549 $\pm$ 0.057 0.523 $\pm$ 0.054 0.557 $\pm$ 0.057
P(True) 0.604 $\pm$ 0.050 0.522 $\pm$ 0.026 0.559 $\pm$ 0.033 0.474 $\pm$ 0.027 0.571 $\pm$ 0.056
LexSim 0.649 $\pm$ 0.055 0.681 $\pm$ 0.046 0.653 $\pm$ 0.053 0.643 $\pm$ 0.054 0.598 $\pm$ 0.060
DegMat 0.734 $\pm$ 0.056 0.672 $\pm$ 0.059 0.616 $\pm$ 0.057 0.608 $\pm$ 0.058 0.678 $\pm$ 0.059
LUQ 0.637 $\pm$ 0.067 0.726 $\pm$ 0.048 0.675 $\pm$ 0.058 0.648 $\pm$ 0.057 0.662 $\pm$ 0.064
KLE 0.333 $\pm$ 0.054 0.341 $\pm$ 0.056 0.357 $\pm$ 0.053 0.360 $\pm$ 0.064 0.338 $\pm$ 0.061
Inv-Entropy 0.788 $\pm$ 0.054 0.740 $\pm$ 0.050 0.703 $\pm$ 0.060 0.780 $\pm$ 0.041 0.695 $\pm$ 0.051
NI-Entropy 0.786 $\pm$ 0.057 0.681 $\pm$ 0.056 0.669 $\pm$ 0.057 0.710 $\pm$ 0.052 0.650 $\pm$ 0.069
NR-Inv-Entropy 0.743 $\pm$ 0.061 0.720 $\pm$ 0.049 0.659 $\pm$ 0.063 0.604 $\pm$ 0.059 0.677 $\pm$ 0.064
WD-px-py 0.518 $\pm$ 0.060 0.303 $\pm$ 0.060 0.443 $\pm$ 0.056 0.573 $\pm$ 0.061 0.605 $\pm$ 0.069
MAX-py-x 0.723 $\pm$ 0.054 0.674 $\pm$ 0.054 0.608 $\pm$ 0.053 0.585 $\pm$ 0.059 0.618 $\pm$ 0.059
PRR ($\uparrow$) Semantic Entropy 0.517 $\pm$ 0.060 0.763 $\pm$ 0.044 0.673 $\pm$ 0.054 0.690 $\pm$ 0.058 0.335 $\pm$ 0.056
VU 0.723 $\pm$ 0.053 0.677 $\pm$ 0.053 0.584 $\pm$ 0.060 0.654 $\pm$ 0.055 0.328 $\pm$ 0.057
P(True) 0.797 $\pm$ 0.042 0.679 $\pm$ 0.050 0.599 $\pm$ 0.049 0.671 $\pm$ 0.041 0.303 $\pm$ 0.056
LexSim 0.810 $\pm$ 0.045 0.770 $\pm$ 0.051 0.654 $\pm$ 0.062 0.767 $\pm$ 0.053 0.356 $\pm$ 0.076
DegMat 0.882 $\pm$ 0.041 0.802 $\pm$ 0.046 0.644 $\pm$ 0.066 0.771 $\pm$ 0.058 0.462 $\pm$ 0.091
LUQ 0.854 $\pm$ 0.043 0.840 $\pm$ 0.045 0.678 $\pm$ 0.066 0.787 $\pm$ 0.052 0.504 $\pm$ 0.094
KLE 0.704 $\pm$ 0.048 0.592 $\pm$ 0.059 0.495 $\pm$ 0.056 0.612 $\pm$ 0.061 0.224 $\pm$ 0.043
Inv-Entropy 0.885 $\pm$ 0.044 0.853 $\pm$ 0.042 0.764 $\pm$ 0.064 0.898 $\pm$ 0.030 0.521 $\pm$ 0.094
NI-Entropy 0.883 $\pm$ 0.043 0.781 $\pm$ 0.053 0.689 $\pm$ 0.071 0.823 $\pm$ 0.055 0.501 $\pm$ 0.098
NR-Inv-Entropy 0.840 $\pm$ 0.054 0.844 $\pm$ 0.045 0.709 $\pm$ 0.068 0.743 $\pm$ 0.064 0.518 $\pm$ 0.087
WD-px-py 0.763 $\pm$ 0.051 0.587 $\pm$ 0.056 0.575 $\pm$ 0.060 0.777 $\pm$ 0.054 0.420 $\pm$ 0.085
MAX-py-x 0.875 $\pm$ 0.038 0.821 $\pm$ 0.048 0.712 $\pm$ 0.055 0.749 $\pm$ 0.062 0.413 $\pm$ 0.081
Brier ($\downarrow$) Semantic Entropy 0.166 $\pm$ 0.023 0.173 $\pm$ 0.020 0.208 $\pm$ 0.017 0.208 $\pm$ 0.020 0.188 $\pm$ 0.018
VU 0.160 $\pm$ 0.022 0.196 $\pm$ 0.017 0.234 $\pm$ 0.010 0.219 $\pm$ 0.017 0.188 $\pm$ 0.020
P(True) 0.172 $\pm$ 0.022 0.215 $\pm$ 0.017 0.233 $\pm$ 0.012 0.215 $\pm$ 0.015 0.189 $\pm$ 0.021
LexSim 0.151 $\pm$ 0.024 0.179 $\pm$ 0.020 0.202 $\pm$ 0.016 0.187 $\pm$ 0.020 0.174 $\pm$ 0.019
DegMat 0.140 $\pm$ 0.021 0.164 $\pm$ 0.018 0.213 $\pm$ 0.014 0.191 $\pm$ 0.018 0.156 $\pm$ 0.019
LUQ 0.148 $\pm$ 0.020 0.159 $\pm$ 0.016 0.191 $\pm$ 0.019 0.180 $\pm$ 0.019 0.151 $\pm$ 0.019
KLE 0.188 $\pm$ 0.021 0.218 $\pm$ 0.016 0.241 $\pm$ 0.006 0.213 $\pm$ 0.018 0.193 $\pm$ 0.021
Inv-Entropy 0.128 $\pm$ 0.020 0.157 $\pm$ 0.018 0.182 $\pm$ 0.020 0.147 $\pm$ 0.017 0.152 $\pm$ 0.020
NI-Entropy 0.124 $\pm$ 0.020 0.164 $\pm$ 0.017 0.196 $\pm$ 0.018 0.168 $\pm$ 0.021 0.156 $\pm$ 0.022
NR-Inv-Entropy 0.138 $\pm$ 0.021 0.159 $\pm$ 0.015 0.199 $\pm$ 0.019 0.188 $\pm$ 0.021 0.165 $\pm$ 0.021
WD-px-py 0.184 $\pm$ 0.019 0.212 $\pm$ 0.016 0.232 $\pm$ 0.009 0.188 $\pm$ 0.018 0.169 $\pm$ 0.021
MAX-py-x 0.148 $\pm$ 0.019 0.177 $\pm$ 0.017 0.207 $\pm$ 0.012 0.189 $\pm$ 0.020 0.169 $\pm$ 0.018
@@
We conducted experiments using two language models: GPT-3.5-Turbo, a black-box model accessed via API, and LLaMA-3.1-8B-Instruct, a grey-box model. We evaluated our framework on datasets spanning three categories: question answering (TriviaQA (Joshi et al., 2017), SciQ (Welbl et al., 2017), Natural Questions (Kwiatkowski et al., 2019)), multiple choice (MMLU (Hendrycks et al., 2020)), and mathematical reasoning (GSM8K (Cobbe et al., 2021)).
@@
We compared our method with various state-of-the-art benchmarks highlighted in Sec. 1, Appendix A and a recent paper (Vashurin et al., 2024) identifying them as top-performers. These include: Semantic Entropy (Farquhar et al., 2024), Verbalized Uncertainty (VU)(Tian et al., 2023), P(True)(Kadavath et al., 2022), Lexical Similarity (LexSim)(Fomicheva et al., 2020), Degree Matrix (DegMat)(Lin et al., 2024), Long-text Uncertainty Quantification (LUQ)(Zhang et al., 2024a), Kernel Language Entropy (KLE)(Nikitin et al., 2024). We also include additional UQ measures based on our framework: (i) NI-Entropy: Non-inverse entropy which uses $P(Y\mid X)$ derived in Sec. 2.3 instead of $P(X\mid Y)$; the rest remains the same. (ii) NR-Inv-Entropy: entropy in (5) without replications. (iii) WD-px-py: Wasserstein distance $\mathsf{WD}(P(X),P(Y))$ ; (iv) MAX-py-x: $\max_{i}P(y_{i}|x_{i})$ from (3).
@@
We evaluate benchmarks using four metrics that are grouped into calibration and uncertainty-based categories. Calibration-based metrics, AUROC, PRR (Malinin and Gales, 2021), and Brier Score (Brier, 1950), assess how well confidence aligns with correctness. The correctness is calculated in the following way. On MMLU, we use exact match of results. On other datasets, we use GPT-3.5-Turbo to judge semantic equivalence. Confidence is defined as the negative of the uncertainty measure for AUROC and PRR, while for Brier Score, we apply isotonic normalization (Zadrozny and Elkan, 2002) to map uncertainty measures to the [0,1] range for all methods. Using correctness as a proxy for UQ has its limitations in open-ended applications or weakly supervised settings where correctness information is not available or not well defined. To circumvent the dependence on correctness, we introduce a new evaluation metric called Temperature Sensitivity of Uncertainty (TSU), a metric that quantifies the proportion of instances where uncertainty increases with temperature. TSU is inspired by the property that LLMs generate responses using a temperature-scaled softmax function, which controls the randomness of token sampling. Higher temperatures flatten token distribution, increasing randomness and uncertainty, while lower temperatures make the sampling process more deterministic (Hinton et al., 2015). Formally, given a sequence of temperature values $\mathbb{T}_{1}<\mathbb{T}_{2}<\dots<\mathbb{T}_{n}$, TSU is defined as:
@@
where $\mathcal{D}$ is the dataset, $S_{x}$ is a question in this dataset, $\texttt{UQ}(S_{x},\mathbb{T})$ represents a UQ subroutine (such as Inv-Entropy) for input $S_{x}$ at temperature $\mathbb{T}$, and $\mathbb{I}(\cdot)$ is the indicator function, which equals 1 if the condition holds and 0 otherwise. A salient feature in the definition of TSU in (7) is that it only depends on $S_{x}$, thus is agnostic to the “ground truth” output $y$. In addition, TSU extends beyond conventional calibration metrics by evaluating the granularity of uncertainty estimation. By leveraging temperature scaling as a probing mechanism, TSU assesses how effectively a method distinguishes between gradations of uncertainty.

### v2

@@ 2.4 Inv-Entropy via bootstrapping and Monte Carlo
A natural measure of uncertainty is the conditional sample entropy $H(X_{n}\mid Y_{n})$, which connects the similarity of the input set $X_{n}$ to the similarity of the corresponding output set $Y_{n}$:
$H(X_{n}\mid Y_{n})\triangleq-\,\operatorname{trace}\!\big(\text{P}_{y}\text{P}_{x}\odot\log(\text{P}_{y}\text{P}_{x})\big)=-\sum_{i=0}^{n}P(x_{i}\mid y_{i})\,\log P(x_{i}\mid y_{i}),$ (5)
where $\odot$ denotes the Hadamard product. This quantity captures the degree of alignment between $\text{P}_{y}$ and $\text{P}_{x}$. High entropy indicates that semantically similar inputs yield divergent outputs, or that semantically distinct outputs correspond to similar inputs, both revealing uncertainty in the model’s behavior. Note that Inv-Entropy represents an unnormalized entropy measure, designed to capture not only the dispersion of $P(x_{i}\mid y_{i})$ but also its magnitude, thereby jointly modeling these two sources of uncertainty.
That said, it is important to realize that $X_{n}$ does not map to a unique $Y_{n}$ due to the inherent stochasticity of $f$ as LLMs can produce different outputs for the same input. Thus, beyond perturbations that capture epistemic uncertainty, replications can help account for aleatoric uncertainty arising from sampling randomness. To this end, each perturbed input $S_{x_{i}}\in\mathcal{P}\mathrm{er}(S_{x_{0}})$ can be queried $r$ times, yielding a replicated output set $\mathcal{R}_{i}=\{y_{i,1},\dots,y_{i,r}\}$ for each question $x_{i}$, where $y_{i,\cdot}=\psi(f(S_{x_{i}}))$.
@@
With this setup, we are able to generate $B$ bootstrap samples $Y_{n}^{(b)}=\{y_{0}^{(b)},\dots,y_{n}^{(b)}\}$ for $b\in\{1,\dots,B\}$, each corresponding to $X_{n}$, where $y_{i}^{(b)}\sim\mathcal{R}_{i}$ is drawn with replacement for $i\in\{0,\dots,n\}$, yielding a transition matrix $\text{P}_{y}^{(b)}$. Each bootstrapped output embedding set $Y_{n}^{(b)}$, together with the input embeddings $X_{n}$, defines a probabilistic instance of our uncertainty measure, Inv-Entropy. The final UQ estimate is obtained by averaging Inv-Entropy over the $B$ bootstrap replicates:
$\widehat{H}(X\mid Y)\triangleq\frac{1}{B}\sum_{b=1}^{B}H(X_{n}\mid Y_{n}^{(b)}).$ (6)
@@
3:LLM Generation: Input each question in $\mathcal{P}er(S_{x_{0}})$, $r$ times into $f$ and obtain $\{\mathcal{R}_{i}\}_{i=0}^{n}$.
4:for $b=1$ to $B$ do
5:  Sample $Y_{n}^{(b)}=\{y_{0}^{(b)},\dots,y_{n}^{(b)}\}$
6:  Compute $\text{P}_{y}^{(b)}$ using (2)
7:  Compute $P^{(b)}(x_{i}|y_{i})=(\text{P}^{(b)}_{y}\cdot\text{P}_{x})[i,i]$ (from (4)) for $\forall$ $i$
8:  Compute $H(X_{n}\mid Y_{n}^{(b)})$ using (5)
@@
2.5 Insights into Inv-Entropy
To provide some insights into our UQ measure, we introduce two parameters: $\epsilon_{x}\in(0,1]$, which controls the input perturbation level, and $\epsilon_{y}\in[0,1]$, which controls the output dispersion level (defined via $a_{\text{Similarity}}$ below). The corresponding transition matrices are then readily derived as:
$a_{\text{Similarity}}(z_{i},z_{j})=\begin{cases}1,&i=j,\\[3.0pt] 1-\epsilon_{z},&i\neq j,\end{cases}\,\rightarrow\,\text{P}_{z}(\epsilon_{z})[i,j]=\begin{cases}\dfrac{1}{(n+1)-n\epsilon_{z}},&i=j,\\[6.0pt] \dfrac{1-\epsilon_{z}}{(n+1)-n\epsilon_{z}},&i\neq j,\end{cases}$
where $z\in\{x,y\}$. With this, the joint conditional probability can be written as: $P(x_{i}\mid y_{i};\epsilon_{x},\epsilon_{y})=\text{P}_{y}(\epsilon_{y})\text{P}_{x}(\epsilon_{x})[i,i]=\frac{1+n-n\epsilon_{x}-n\epsilon_{y}+n\epsilon_{x}\epsilon_{y}}{(n+1-n\epsilon_{x})(n+1-n\epsilon_{y})}.$ We can then show that for the same input perturbation level $\epsilon_{x}$, a larger output dispersion $\epsilon_{y}$ corresponds to higher uncertainty, consistent with the definition of Inv-Entropy. To see this, notice that the derivative of $P(x_{i}\mid y_{i};\epsilon_{x},\epsilon_{y})$ with respect to $\epsilon_{y}$ is $\frac{\partial P(x_{i}\mid y_{i};\epsilon_{x},\epsilon_{y})}{\partial\epsilon_{y}}=\frac{n\epsilon_{x}}{(n+1-n\epsilon_{x})(n+1-n\epsilon_{y})^{2}}>0.$ Now, if we assume all $P(x_{i}\mid y_{i};\epsilon_{x},\epsilon_{y})$ are smaller than $\frac{1}{e}$ (a condition generally satisfied for large $n$), then the Inv-Entropy function $H$ in (5) becomes also increasing in $P(x_{i}\mid y_{i};\epsilon_{x},\epsilon_{y})$ and hence in $\epsilon_{y}$.
@@ 2.6 GAAP
In this section, we present a genetic algorithm-based adversarial perturbation (GAAP) that progressively modifies the semantic input $S_{x_{0}}$ to generate controlled perturbations, $\mathcal{P}er(S_{x_{0}})$. Below we highlight our overarching framework, while algorithmic details are relegated to Appendix C. As shown in Fig. 4, the process consists of an initialization step and multiple iterative procedures. In the initialization step, we construct a population $\text{Pop}_{0}(S_{x_{0}})$ consisting of perturbed versions of $S_{x_{0}}$. More specifically, each text in $\text{Pop}_{0}(S_{x_{0}})$ is derived from $S_{x_{0}}$ by replacing one keyword with a synonym, hypernym, hyponym from WordNet (10.1145/219717.219748), or a deletion.
Figure 4: Illustration of GAAP on a TriviaQA (joshi2017triviaqa) question.
@@ 3 Experiments
AUROC ($\uparrow$) Semantic Entropy 0.579 $\pm$ 0.044 0.679 $\pm$ 0.045 0.521 $\pm$ 0.034 0.518 $\pm$ 0.048 0.589 $\pm$ 0.052
Kernel Entropy 0.687 $\pm$ 0.062 0.685 $\pm$ 0.063 0.556 $\pm$ 0.055 0.653 $\pm$ 0.059 0.560 $\pm$ 0.060
VU 0.695 $\pm$ 0.060 0.480 $\pm$ 0.060 0.533 $\pm$ 0.056 0.523 $\pm$ 0.054 0.557 $\pm$ 0.057
P(True) 0.604 $\pm$ 0.050 0.522 $\pm$ 0.026 0.519 $\pm$ 0.020 0.474 $\pm$ 0.027 0.571 $\pm$ 0.056
LexSim 0.649 $\pm$ 0.055 0.681 $\pm$ 0.046 0.518 $\pm$ 0.055 0.643 $\pm$ 0.054 0.598 $\pm$ 0.060
DegMat 0.734 $\pm$ 0.056 0.672 $\pm$ 0.059 0.551 $\pm$ 0.052 0.608 $\pm$ 0.058 0.678 $\pm$ 0.059
LUQ 0.637 $\pm$ 0.067 0.726 $\pm$ 0.048 0.627 $\pm$ 0.055 0.648 $\pm$ 0.057 0.662 $\pm$ 0.064
KLE 0.333 $\pm$ 0.054 0.341 $\pm$ 0.056 0.410 $\pm$ 0.060 0.360 $\pm$ 0.064 0.338 $\pm$ 0.061
Inv-Entropy 0.788 $\pm$ 0.054 0.740 $\pm$ 0.050 0.661 $\pm$ 0.052 0.780 $\pm$ 0.041 0.695 $\pm$ 0.051
NI-Entropy 0.786 $\pm$ 0.057 0.681 $\pm$ 0.056 0.637 $\pm$ 0.053 0.710 $\pm$ 0.052 0.650 $\pm$ 0.069
NR-Inv-Entropy 0.743 $\pm$ 0.061 0.720 $\pm$ 0.049 0.627 $\pm$ 0.054 0.604 $\pm$ 0.059 0.677 $\pm$ 0.064
WD-px-py 0.518 $\pm$ 0.060 0.303 $\pm$ 0.060 0.558 $\pm$ 0.055 0.573 $\pm$ 0.061 0.605 $\pm$ 0.069
MAX-py-x 0.723 $\pm$ 0.054 0.674 $\pm$ 0.054 0.547 $\pm$ 0.051 0.585 $\pm$ 0.059 0.618 $\pm$ 0.059
PRR ($\uparrow$) Semantic Entropy 0.517 $\pm$ 0.060 0.763 $\pm$ 0.044 0.505 $\pm$ 0.049 0.690 $\pm$ 0.058 0.335 $\pm$ 0.056
Kernel Entropy 0.794 $\pm$ 0.052 0.812 $\pm$ 0.039 0.573 $\pm$ 0.068 0.768 $\pm$ 0.057 0.333 $\pm$ 0.054
VU 0.723 $\pm$ 0.053 0.677 $\pm$ 0.053 0.537 $\pm$ 0.053 0.654 $\pm$ 0.055 0.328 $\pm$ 0.057
P(True) 0.797 $\pm$ 0.042 0.679 $\pm$ 0.050 0.502 $\pm$ 0.050 0.671 $\pm$ 0.041 0.303 $\pm$ 0.056
LexSim 0.810 $\pm$ 0.045 0.770 $\pm$ 0.051 0.563 $\pm$ 0.064 0.767 $\pm$ 0.053 0.356 $\pm$ 0.076
DegMat 0.882 $\pm$ 0.041 0.802 $\pm$ 0.046 0.549 $\pm$ 0.069 0.771 $\pm$ 0.058 0.462 $\pm$ 0.091
LUQ 0.854 $\pm$ 0.043 0.840 $\pm$ 0.045 0.595 $\pm$ 0.066 0.787 $\pm$ 0.052 0.504 $\pm$ 0.094
KLE 0.704 $\pm$ 0.048 0.592 $\pm$ 0.059 0.449 $\pm$ 0.059 0.612 $\pm$ 0.061 0.224 $\pm$ 0.043
Inv-Entropy 0.885 $\pm$ 0.044 0.853 $\pm$ 0.042 0.614 $\pm$ 0.067 0.898 $\pm$ 0.030 0.521 $\pm$ 0.094
NI-Entropy 0.883 $\pm$ 0.043 0.781 $\pm$ 0.053 0.592 $\pm$ 0.064 0.823 $\pm$ 0.055 0.501 $\pm$ 0.098
NR-Inv-Entropy 0.840 $\pm$ 0.054 0.844 $\pm$ 0.045 0.576 $\pm$ 0.069 0.743 $\pm$ 0.064 0.518 $\pm$ 0.087
WD-px-py 0.763 $\pm$ 0.051 0.587 $\pm$ 0.056 0.586 $\pm$ 0.065 0.777 $\pm$ 0.054 0.420 $\pm$ 0.085
MAX-py-x 0.875 $\pm$ 0.038 0.821 $\pm$ 0.048 0.536 $\pm$ 0.066 0.749 $\pm$ 0.062 0.413 $\pm$ 0.081
Brier ($\downarrow$) Semantic Entropy 0.166 $\pm$ 0.023 0.173 $\pm$ 0.020 0.242 $\pm$ 0.006 0.208 $\pm$ 0.020 0.188 $\pm$ 0.018
Kernel Entropy 0.160 $\pm$ 0.025 0.153 $\pm$ 0.022 0.221 $\pm$ 0.011 0.179 $\pm$ 0.018 0.190 $\pm$ 0.017
VU 0.160 $\pm$ 0.022 0.196 $\pm$ 0.017 0.223 $\pm$ 0.014 0.219 $\pm$ 0.017 0.188 $\pm$ 0.020
P(True) 0.172 $\pm$ 0.022 0.215 $\pm$ 0.017 0.244 $\pm$ 0.005 0.215 $\pm$ 0.015 0.189 $\pm$ 0.021
LexSim 0.151 $\pm$ 0.024 0.179 $\pm$ 0.020 0.225 $\pm$ 0.010 0.187 $\pm$ 0.020 0.174 $\pm$ 0.019
DegMat 0.140 $\pm$ 0.021 0.164 $\pm$ 0.018 0.229 $\pm$ 0.012 0.191 $\pm$ 0.018 0.156 $\pm$ 0.019
LUQ 0.148 $\pm$ 0.020 0.159 $\pm$ 0.016 0.208 $\pm$ 0.014 0.180 $\pm$ 0.019 0.151 $\pm$ 0.019
KLE 0.188 $\pm$ 0.021 0.218 $\pm$ 0.016 0.244 $\pm$ 0.006 0.213 $\pm$ 0.018 0.193 $\pm$ 0.021
Inv-Entropy 0.128 $\pm$ 0.020 0.157 $\pm$ 0.018 0.201 $\pm$ 0.014 0.147 $\pm$ 0.017 0.152 $\pm$ 0.020
NI-Entropy 0.124 $\pm$ 0.020 0.164 $\pm$ 0.017 0.204 $\pm$ 0.014 0.168 $\pm$ 0.021 0.156 $\pm$ 0.022
NR-Inv-Entropy 0.138 $\pm$ 0.021 0.159 $\pm$ 0.015 0.208 $\pm$ 0.013 0.188 $\pm$ 0.021 0.165 $\pm$ 0.021
WD-px-py 0.184 $\pm$ 0.019 0.212 $\pm$ 0.016 0.225 $\pm$ 0.010 0.188 $\pm$ 0.018 0.169 $\pm$ 0.021
MAX-py-x 0.148 $\pm$ 0.019 0.177 $\pm$ 0.017 0.229 $\pm$ 0.011 0.189 $\pm$ 0.020 0.169 $\pm$ 0.018
@@
We conducted experiments using two language models: GPT-3.5-Turbo, a black-box model accessed via API, and LLaMA-3.1-8B-Instruct, a grey-box model. We evaluated our framework on datasets spanning three categories: question answering (TriviaQA (joshi2017triviaqa), SciQ (welbl-etal-2017-crowdsourcing), Natural Questions (kwiatkowski2019natural) (NQ, long-answer questions with details in Appendix D.3)), multiple choice (MMLU (hendrycks2020measuring)), and mathematical reasoning (GSM8K (cobbe2021training)).
@@
We compared our method with various state-of-the-art benchmarks highlighted in Sec. 1, Appendix A and a recent paper (DBLP:journals/corr/abs-2406-15627) identifying them as top-performers. These include: Semantic Entropy (farquhar2024detecting), Kernel Entropy (gruber2023bias), Verbalized Uncertainty (VU) (tian2023justaskcalibrationstrategies), P(True) (kadavath2022languagemodelsmostlyknow), Lexical Similarity (LexSim) (fomicheva2020unsupervised), Degree Matrix(DegMat) (lin2024generating), Long-text Uncertainty Quantification (LUQ) (zhang-etal-2024-luq), Kernel Language Entropy (KLE) (nikitin2024kernellanguageentropyfinegrained). We also include additional UQ measures based on our framework: (i) NI-Entropy: Non-inverse entropy which uses $P(Y\mid X)$ derived in Sec. 2.3 instead of $P(X\mid Y)$; the rest remains the same. (ii) NR-Inv-Entropy: entropy in (5) without replications. (iii) WD-px-py: Wasserstein distance $\mathsf{WD}(P(X),P(Y))$ ; (iv) MAX-py-x: $\max_{i}P(y_{i}|x_{i})$.
@@
We evaluate performance using four metrics grouped into correctness-based and uncertainty-based categories. Correctness-based metrics: AUROC, PRR (malinin2021uncertainty), and Brier Score (brier1950verification), measure how well confidence aligns with correctness. For MMLU, correctness is defined via exact match, while for other datasets we use GPT-3.5-Turbo to assess whether a generated response is semantically equivalent to the reference (ground-truth) answer. Confidence is taken as the negative of the UQ measure (i.e. Inv-Entropy) for AUROC and PRR. For the Brier Score, we apply isotonic normalization (zadrozny2002transforming) to map uncertainty scores to the [0,1] range. Correctness-based metrics, however, rely on ground truth and may fail in open-ended or weakly supervised settings. To address this, we propose the Temperature Sensitivity of Uncertainty (TSU), which quantifies how often uncertainty increases with temperature. Since higher temperatures flatten the softmax distribution, they should yield greater randomness and uncertainty (hinton2015distilling). Formally, given a sequence of temperature values $\mathbb{T}_{1}<\mathbb{T}_{2}<\dots<\mathbb{T}_{n}$, TSU is defined as:
@@
where $\mathcal{D}$ is the dataset, $S_{x}$ is a question in this dataset, $\texttt{UQ}(S_{x},\mathbb{T})$ represents a UQ subroutine (such as Inv-Entropy) for input $S_{x}$ at temperature $\mathbb{T}$, and $\mathbb{I}(\cdot)$ is the indicator functionA salient feature in the definition of TSU in (7) is that it only depends on $S_{x}$, thus is agnostic to the “ground truth” output $y$. In addition, TSU extends beyond conventional correctness-based metrics by evaluating the granularity of uncertainty estimation. By leveraging temperature scaling as a probing mechanism, TSU assesses how effectively a method distinguishes between gradations of uncertainty.